OpenStudy (anonymous):

The solution set of the equation |x - 1| = 3 is _____

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (fibonaccichick666):

Please show us your work, what have you done so far?

OpenStudy (anonymous):

write it as two equations

OpenStudy (anonymous):

x-1=3 -(x-1)=3

OpenStudy (anonymous):

solve these two

OpenStudy (anonymous):

also thre is one quith method

OpenStudy (uri):

|x - 1| = 3 x-1=-3 x-1=3

OpenStudy (uri):